# Decay rates to equilibrium in a nonlinear subdiffusion equation with two counteracting terms

### v1

@@ 1 Introduction
Here $p$, $q$, $f$ and often also the state $u$ are typically nonnegative, so that the two terms $q(x)u$ and $p(x)f(u)$ play counteracting roles in the PDE.
While spatial inhomogeneity of the coefficients $p$ and $q$ arises quite naturally in the mentioned applications, the literature on the non-constant setting is quite scarce; this is probably due to the fact that most of the interest went into investigation of travelling wave solutions, which are largely inhibited by spatially varying coefficients $p$ and $q$, though. The large time limit has already been studied in [3] and steady state solutions of the Fisher-KPP equation as well as their relations to travelling wave solutions have been subject of intensive research, see, e.g., [4, 5] for the non-local setting. The same holds true for the steady state solutions of the Allen-Cahn equation, that can be characterized as minimizers of an appropriate energy functional, see, e.g., [1, 15]; in particular, we point to [6], [8, Section 6.4.2] for the time-fractional case.
@@
where $\Omega\subseteq\mathbb{R}^{d}$, $d\in\{1,2,3\}$ is a smooth and bounded domain and the operator $B$ defining the boundary conditions is of Dirichlet or Neumann type; we expect the analysis to carry over to more general boundary conditions, such as impedance or mixed conditions.
@@
as this can be achieved without changing the problem by making the replacements $f\to f(u)-f^{\prime}(0)u-f(0)$, $q\to q+f^{\prime}(0)p$, $r\to r-f(0)p$. We will require $p$ and $f$ to be positive, but do not impose any sign conditions on $r$. Essential sign (in fact, monotonicity) conditions that indirectly also affect $q$ will be imposed below, cf. (12), (24) below.
@@
and satisfies (among other properties) the coercivity estimate [10, Lemma 4.18]
@@ 2 Main results
$\|\partial_{t}^{\alpha}u(t)\|_{H^{\tilde{s}}(\Omega)}^{2}\leq C\Psi(t),\quad\|u_{t}(t)\|_{H^{s}(\Omega)}^{2}\leq C\Psi(t)\quad t>0$ (9)
$\tilde{s}\in\{-1,0\}$, $s\in\{0,1,2\}$, cf. Theorems 2.1, 2.2, 2.3.
@@
  • development and use of Gronwall type estimates; alternatively we also employ Tauberian theorems for quantifying the large time asymptotic behaviour.
@@
Results related to (8) are, e.g., [8, Sections 6.4.1 and 6.4.2], [9, Section 4.2.6], on the case $p\equiv 1$, $q=0$, and [13, Section 4.6], where the latter is on non-autonomous but linear problems. However, the current setting is not covered by these results. In particular, the approach from [8, Sections 6.4.1] (see also [9, Section 4.2.6]) of representing the solution via the variation of constants formula
@@
with the Mittag-Leffler function $E_{\alpha,1}$ and an eigensystem $(\lambda_{j},\varphi_{j})_{j\in\mathbb{N}}$ of the elliptic operator, would require higher spatial regularity on the coefficients $p$ and $q$, that cannot be reproduced by the fixed point operator $\mathcal{T}$ on which the reconstruction scheme in [11] relies. It is thus important to be able to work in the coefficient low regularity setting (10).
Remarkable on the forward problem are the counteracting roles of the $p$ and $q$ terms, when having nonnegativity of these two coefficients as well as a nonnegative function $f$ in mind, as typical in applications. Depending on the monotonicity behaviour of $f$ and on which of the terms dominates, this can result in exponential decay or blow-up of solutions or in convergence to a steady state $u^{\infty}$ as time tends to infinity, that has to satisfy the elliptic PDE
@@
Existence of a solution $u^{\infty}\in H^{2}(\Omega)$ holds under mild conditions on the coefficients and the function $f$ from monotonicity arguments, see, e.g., [17, Theorem 4.4].
@@
with $\Psi$ as in (8), $\omega=c_{1}>0$ with $X\in\{H^{-1}(\Omega),\,L^{2}(\Omega)\}$.
@@
but expect our results to remain valid, provided $a(t,x)\to a^{\infty}(x)$, $B\to B^{\infty}$, as $t\to\infty$ sufficiently fast. Correspondingly we use $V=H^{1}(\Omega)$ or $V=H_{0}^{1}(\Omega)$ as function space (and by a slight abuse of notation, write $H^{-1}(\Omega)$ for its topological dual) for $u(t)$.
@@
For $u_{0}\in H^{1}(\Omega)$, $f\in C^{2}(\mathbb{R})$, under conditions (2) (3), (4), (12), (13), $r\in L^{\infty}(0,\infty;X)$, (16), (17), there exist $\rho_{0}>0$, $C>0$, $0<c_{0}<c_{1}$ independent of $T$ such that $u^{\sim}=u-u^{\infty}$ (where $u$ solves (1) and $u^{\infty}$ solves (11)) satisfies the bound
@@
To prove existence, one can follow the standard approch of Galerkin approximation, energy estimates and weak limits (like in, e.g., [8, Section 6.4.2], but using the energy estimates in the proof of Theorem 2.1, that avoid differentiation of $p$ and $q$). Uniqueness follows by the testing that we have employed for obtaining lower order estimates, cf. (26). ∎
@@
    If $\|r_{t}(t)\|_{L^{2}(\Omega)}^{2}\leq C_{r}\Psi(t)$ for some $\omega:=c_{1}>0$, then there exist $C,\,\omega>0$ independent of $T$, $t_{0}$, such that
@@
An improved decay estimate as in [13, Section 4.6] would extend the range of possible $\alpha$ values to $\alpha\in(\frac{1}{2},1)$
@@
and apply Gronwall’s inequality (see, e.g., [8, Theorem 4.2], [10, Lemma A.1.]) to obtain
@@
Note that we do not integrate the term contaning $p^{\infty}$ by parts, since this would leads to a gradient of $p^{\infty}$, but estimate it by
@@
The estimates from Proposition 3.1 allow us to prove decay rates $\|u^{\sim}(t)\|^{2}=O(\Psi(t))$ with $\Psi$ as in (8) as well known in case $\alpha=1$, whereas in case $\alpha\in(0,1)$ this can be deduced from results in, e.g., [13, Chapter 5]. (Note that there even better decay rates are shown; however, without inhomogeneity.) For the convenience of the reader, we summarize this in a lemma and provide its proof in the appendix.
@@
A Tauberian Theorem (e.g., [10, Theorem 3.14] with $dw=\eta_{j}(t)\,dt$), which is also valid for $\alpha=1$, thus yields
@@
The implication (47) $\Rightarrow$ (48) remains valid with the upper bound $1$ replaced by $\frac{1}{2}$, thus implying
@@
We test with $\mu\partial_{t}^{\tilde{\alpha}}u^{\sim}-\mathbb{L}\partial_{t}^{\tilde{\alpha}}u^{\sim}$, thus focusing ton the higher regularity setting $s=1$. This, analogously to (28), yields
@@
with
@@
In case $\tilde{\alpha}<1$, the convolution term on the right hand side of (50) is problematic, since it contains $u^{\sim}_{t}$, which is a higher derivative than what we can estimate in case $\tilde{\alpha}<1$. In order to tackle it, we would need a commutator estimate. However such an estimate typically takes the form [14, Corollary 1.4 (2) with $A^{\tilde{\alpha}}=\partial_{t}^{\tilde{\alpha}}$]
@@
for $0\leq\alpha_{1}\leq\tilde{\alpha}$ and the nonlocal character of the right hand side does not allow to reproduce decay as $t\to\infty$. On the other hand, the fractional Leibniz rule
$\partial_{t}^{\tilde{\alpha}}(h\,(g-g(0)))=\sum_{k=0}^{\infty}\binom{\tilde{\alpha}}{k}\,h^{(k)}\,{\partial}_{t}^{\tilde{\alpha}-k}(g-g(0))$
involves arbitrarily high derivatives of the multiplier $h$.
@@
where we can use the previously established decay estimate (8) with $s=1$ on $u^{\sim}(t)$ to deduce existence of $t_{0}\leq 0$ such that
@@
$\leq\frac{\min\{\mu\,\min\{\tfrac{\underline{c}}{2},c_{D}\},\,1\}}{4(C_{\text{ell}}^{\Omega})^{2}}\qquad t\geq t_{0}.$
@@
First of all, note that the regularity assumption on $\eta$ imply continuity of $\eta$ and $\partial_{t}^{\alpha}\eta$ on $(0,T)$ and therefore the word “almost” can be skipped in (36). Since $\nu$ satisfies (36) with equality (cf., e.g. [10, Theorem 5.4]) for all $t\in(0,T)$, as well as $\nu(0)=\eta(0)$, we obtain, for the difference $\eta-\nu$,
@@
Lemma ([13, Lemma 5.2]) Let $f\in W^{1,1}(0,T)$ attain its maximum over the interval $[0,T]$ at a point $t_{0}\in(0,T]$. If for each $\underline{t}\in(0,T)$, there exists $\beta\in(0,1]$ such that $f\in W^{1,1/(1-\beta)}(\underline{t},T)$, then $(\partial_{t}^{\alpha}f)(t_{0})\geq 0$ for every $\alpha\in(0,\beta)$.
@@
In case $\phi(t)=\phi_{0}t^{-\alpha}$ we can estimate similarly to the proof of [10, Lemma 10.11] make use of the asympotics of Mittag Leffler functions and their interrelations, cf., e.g., [10, Section 3.4], more precisely

### v2

@@ 1 Introduction
Here $p$, $q$, and often also $f^{\prime}$ are typically nonnegative, so that the two terms $q(x)u$ and $p(x)f(u)$ play counteracting roles in the PDE.
While spatial inhomogeneity of the coefficients $p$ and $q$ arises quite naturally in the mentioned applications, the literature on the non-constant setting is quite scarce; this is probably due to the fact that most of the interest went into investigation of travelling wave solutions, which are largely inhibited by spatially varying coefficients $p$ and $q$, though. The large time limit has already been studied in [4] and steady state solutions of the Fisher-KPP equation as well as their relations to travelling wave solutions have been subject of intensive research, see, e.g., [5, 6] for the non-local setting. The same holds true for the steady state solutions of the Allen-Cahn equation, that can be characterized as minimizers of an appropriate energy functional, see, e.g., [1, 16]; in particular, we point to [7], [9, Section 6.4.2] for the time-fractional case.
@@
where $\Omega\subseteq\mathbb{R}^{d}$, $d\in\{1,2,3\}$ is a smooth and bounded domain and the operator $B$ defining the boundary conditions is linear.
@@
as this can be achieved without changing the problem by making the replacements $f\to f(u)-f^{\prime}(0)u-f(0)$, $q\to q+f^{\prime}(0)p$, $r\to r-f(0)p$. We will require $p$ and $f^{\prime}$ to be positive, but do not impose any sign conditions on $r$. Essential sign (in fact, monotonicity) conditions that indirectly also affect $q$ will be imposed below, cf. (12), (24) below.
@@
and satisfies (among other properties) the coercivity estimate [2, Lemma 1]
@@ 2 Main results
$\|\partial_{t}^{\alpha}u(t)\|_{H^{s}(\Omega)}^{2}\leq C\Psi(t),\quad\|u_{t}(t)\|_{H^{s}(\Omega)}^{2}\leq C\Psi(t)\quad t>0$ (9)
$s\in\{-1,0,1,2\}$, cf. Theorems 2.1, 2.2, 2.3, Remark 2.1 under various different assumptions.
@@
  • use (and partly also development) of Gronwall type estimates; alternatively we also employ Tauberian theorems for quantifying the large time asymptotic behaviour.
@@
Results related to (8) are, e.g., [9, Sections 6.4.1 and 6.4.2], [10, Section 4.2.6], on the case $p\equiv 1$, $q=0$, and [14, Section 4.6], where the latter is on non-autonomous but linear problems. However, the current setting is not covered by these results. In particular, the approach from [9, Sections 6.4.1] (see also [10, Section 4.2.6]) of representing the solution via the variation of constants formula
@@
with the Mittag-Leffler function $E_{\alpha,1}$ and an eigensystem $(\lambda_{j},\varphi_{j})_{j\in\mathbb{N}}$ of the elliptic operator, would require higher spatial regularity on the coefficients $p$ and $q$, that cannot be reproduced by the fixed point operator $\mathcal{T}$ on which the reconstruction scheme in [12] relies. It is thus important to be able to work in the low regularity coefficient setting (10).
Remarkable on the forward problem are the counteracting roles of the $p$ and $q$ terms, when having nonnegativity of these two coefficients as well as a monotonically increasing function $f$ in mind, as typical in applications. Depending on the monotonicity behaviour of $f$ and on which of the terms dominates, this can result in exponential decay or blow-up of solutions or in convergence to a steady state $u^{\infty}$ as time tends to infinity, that has to satisfy the elliptic PDE
@@
Existence of a solution $u^{\infty}\in H^{2}(\Omega)$ holds under mild conditions on the coefficients and the function $f$ from monotonicity arguments, see, e.g., [18, Theorem 4.4].
@@
with $\Psi$ as in (8), $\omega=\omega_{r}>0$ and $X\in\{H^{-1}(\Omega),\,L^{2}(\Omega)\}$.
@@
but expect our results to remain valid, provided $a(t,x)\to a^{\infty}(x)$, $B\to B^{\infty}$, as $t\to\infty$ sufficiently fast. We also expect the analysis to carry over to more general boundary conditions, such as impedance or mixed conditions.
Correspondingly we use $V=H^{1}(\Omega)$ or $V=H_{0}^{1}(\Omega)$ as function space (and by a slight abuse of notation, write $H^{-1}(\Omega)$ for its topological dual) for $u(t)$.
@@
For $u_{0}\in H^{1}(\Omega)$, $f\in C^{2}(\mathbb{R})$, under conditions (2) (3), (4), (12), (13), $r\in L^{\infty}(0,\infty;X)$, (16), (17), there exist $\rho_{0}>0$, $C>0$, independent of $T$ such that $u^{\sim}=u-u^{\infty}$ (where $u$ solves (1) and $u^{\infty}$ solves (11)) satisfies the bound
@@
To prove existence, one can follow the standard approch of Galerkin approximation, energy estimates and weak limits (like in, e.g., [9, Section 6.4.2], but using the energy estimates in the proof of Theorem 2.1, that avoid differentiation of $p$ and $q$). Uniqueness follows by the testing that we employ for obtaining lower order estimates, cf. (26). ∎
@@
    If $\|r_{t}(t)\|_{L^{2}(\Omega)}^{2}\leq C_{r}\Psi(t)$ for some $\omega:=\omega_{r}>0$, then there exist $C,\,\omega>0$ independent of $T$, $t_{0}$, such that
@@
An improved decay estimate as in [14, Section 4.6] would extend the range of possible $\alpha$ values to $\alpha\in(\frac{1}{2},1)$
@@
and apply Gronwall’s inequality (see, e.g., [9, Theorem 4.2], [11, Lemma A.1.]) to obtain
@@ 3.1 Energy estimates
Note that we do not integrate the term contaning $p^{\infty}$ by parts, since this would lead to a gradient of $p^{\infty}$, but estimate it by
@@
3.2 Gronwall type estimates
The estimates from Proposition 3.1 allow us to prove decay rates $\|u^{\sim}(t)\|^{2}=O(\Psi(t))$ with $\Psi$ as in (8) as well known in case $\alpha=1$, whereas in case $\alpha\in(0,1)$ this can be deduced from results in, e.g., [14, Chapter 5]. (Note that there even better decay rates are shown; however, without inhomogeneity.) For the convenience of the reader, we summarize this in a lemma and provide its proof in the appendix.
@@ 3.2 Gronwall type estimates
A Tauberian Theorem (e.g., [11, Theorem 3.14] with $dw=\eta_{j}(t)\,dt$), which is also valid for $\alpha=1$, thus yields
@@
$c\,\|u^{\sim}(t)\|_{H^{s+1}(\Omega)}^{2}\bigl(1-D^{s}(t)\bigr)$
for $s\in\{0,1\}$, and set $c_{0}=c_{1}=\frac{c}{2}$,
The implication (47) $\Rightarrow$ (48) remains valid with the upper bound 11 replaced by $\frac{1}{2}$, thus implying
@@
We test with $\mu\partial_{t}^{\tilde{\alpha}}u^{\sim}-\mathbb{L}\partial_{t}^{\tilde{\alpha}}u^{\sim}$, thus focusing on the higher regularity setting $s=1$. This, analogously to (28), yields
@@
with the indicator variable $\chi_{A}=\begin{cases}1&\text{ if $A$ holds}\\ 0&\text{ else}\end{cases}$ and
@@
In case $\tilde{\alpha}<1$, the convolution term on the right hand side of (50) is problematic, since it contains $u^{\sim}_{t}$, which is a higher derivative than what we can estimate in case $\tilde{\alpha}<1$. In order to tackle it, we would need a commutator estimate. However such an estimate typically takes the form [15, Corollary 1.4 (2) with $A^{\tilde{\alpha}}=\partial_{t}^{\tilde{\alpha}}$]
@@
for $0\leq\alpha_{1}\leq\tilde{\alpha}$ and the nonlocal character of the right hand side does not allow to reproduce decay as $t\to\infty$.
@@
where we can use the previously established decay estimate (8) with $s=1$ on $u^{\sim}(t)$ to deduce existence of $t_{1}\geq 0$ such that
@@
$\leq\frac{\min\{\mu\,\min\{\tfrac{\underline{c}}{2},c_{D}\},\,1\}}{4(C_{\text{ell}}^{\Omega})^{2}}\qquad t\geq t_{1}.$
@@
First of all, note that the regularity assumption on $\eta$ imply continuity of $\eta$ and $\partial_{t}^{\alpha}\eta$ on $(0,T)$ and therefore the word “almost” can be skipped in (36). Since $\nu$ satisfies (36) with equality (cf., e.g. [11, Theorem 5.4]) for all $t\in(0,T)$, as well as $\nu(0)=\eta(0)$, we obtain, for the difference $\eta-\nu$,
@@
Lemma ([14, Lemma 5.2]) Let $f\in W^{1,1}(0,T)$ attain its maximum over the interval $[0,T]$ at a point $t_{0}\in(0,T]$. If for each $\underline{t}\in(0,T)$, there exists $\beta\in(0,1]$ such that $f\in W^{1,1/(1-\beta)}(\underline{t},T)$, then $(\partial_{t}^{\alpha}f)(t_{0})\geq 0$ for every $\alpha\in(0,\beta)$.
@@
In case $\phi(t)=\phi_{0}t^{-\alpha}$ we can estimate similarly to the proof of [11, Lemma 10.11] make use of the asympotics of Mittag Leffler functions and their interrelations, cf., e.g., [11, Section 3.4], more precisely